Step by step
Problem No. 4. 10 pts Given Histogram shows the distribution of serum cholesterol level in milligrams per deciliter for a sample of men. Healthy serum cholesterol level is less than 200 mg/dL. Borderline high risk level: 200- 240 mg/dL. High risk level: greater than 240 mg/dL. 0 20 0.15 0.10 0.05 0.00 100 200 Serum choiesterol (mgidl Use the histogram to estimate the following: (a) Is the percentage of men with cholesterol levels below 240 mg/dL closest to 30%, 50% or 70%? (b) In which interval are there more men: 100-200 mg/dL or 240-340 mg/dL.? Solution
Answer #1

From the given histogram, the frequency table obtained is as follows

Class Interval Frequency Cumulative Frequency
100-120 2 2
120-140 4 6
140-160 3 9
160-180 12 21
180-200 13 34
200-220 19 53
220-240 15 68
240-260 14 82
260-280 8 90
280-300 5 95
300-320 1 96
320-340 4 100

a) The percentage of men with cholesterol level below 240mg/dL is closest to 70%. i.e., the cumulative frequency below 240 mg/dL is 68 (2+4+3+12+13+19+15)

b) between the interval 100-200 the number of men = 34

between the interval 240 -340 the number of men = 32 (14+8+5+1+4)

Hence we can say that there are more men between the interval 100-200 as compared to the interval 240-340